OpenStudy (anonymous):

(8y – 7) – (5y + 1) = 4.

OpenStudy (anonymous):

3y-8=4 3y=4+8 3y=12 3y/3=12/3 y=4

OpenStudy (anonymous):

8y-7 -5y-1=4 3y-8=4 +8 +8 3y=12 divide both sides by 3 y=4